# A Survey on Dynamic Neural Networks: from Computer Vision to Multi-modal Sensor Fusion

### v3

@@ Abstract
Model compression is essential in the deployment of large Computer Vision models on embedded devices. However, static optimization techniques (e.g. pruning, quantization, etc.) neglect the fact that different inputs have different complexities, thus requiring different amount of computations. Dynamic Neural Networks allow to condition the number of computations to the specific input. The current literature on the topic is very extensive and fragmented. We present a comprehensive survey that synthesizes and unifies existing Dynamic Neural Networks research in the context of Computer Vision. Additionally, we provide a logical taxonomy based on which component of the network is adaptive: the output, the computation graph or the input. Furthermore, we argue that Dynamic Neural Networks are particularly beneficial in the context of Sensor Fusion for better adaptivity, noise reduction and information prioritization. We present preliminary works in this direction. We complement this survey with a curated repository listing all the surveyed papers, each with a brief summary of the solution and the code base when available: https://github.com/DTU-PAS/awesome-dynn-for-cv.
@@ 1 Introduction
Furthermore, we argue that the adaptability of Dynamic Neural Networks is a particularly beneficial property in the context of Sensor Fusion—however, this is currently an underexplored research area. Studies suggest that, in the human brain, distinct parts are highly specialized in different tasks (Kandel et al., ), e.g. spatial information and object information are processed separately (Goodale and Milner, 1992). Furthermore, studies on different living organisms have shown how the loss of a sense gets compensated from the remaining sensory modalities (Rabinowitch et al., 2016). In Computer Vision, Multi-modal Sensor Fusion can be defined as the aptitude of a model to combine, interpret, and reason on information coming from different sensors and express various aspects of the same environment, aiming at a better perception of the surroundings in different contexts and conditions (Fung et al., 2017).
We observe that Sensor Fusion models can strongly benefit from the use of adaptive techniques for multiple reasons. Firstly, they allow sample-aware model efficiency. Furthermore, the learned adaptivity produces robustness against the noise generated by non-optimal environmental conditions in which the sensors are exposed. Finally, there is an intrinsic ability of these techniques to identify critical information by design. This allows to organize the processing according to input importance and computation budget constraints. As can be seen in Figure 2, Dynamic Neural Networks are gaining traction, with the number of papers steadily increasing year by year. While—as will be discussed later in this section—existing survey papers have covered similar directions, there is still a lack of a survey paper and a clear taxonomy that focuses on Computer Vision.
Figure 2: Overview of the publications considered in this survey, grouped by year and topic. In total, 163 publications have been reviewed: 62 for the Early Exits Section, 45 for the Dynamic Routing Section, 27 for the Token Skimming Section, and 16 in the Dynamic Sensor Fusion Section.
@@ 2.1 Early exits
The idea at the base of the development of Early Exits networks is rooted in the assumption that samples in a dataset can vary a lot in complexity. Following this principle, traditional Neural Networks tend to be overparametrized for most of the samples, which could potentially be inferred with just a fraction of the parameters used. Furthermore, in the case of easy samples, Convolutional Neural Networks (CNNs) have been shown to be capable of extracting features in the early layers that are already sufficient for a confident classification (Panda et al., 2016). For these reasons, Early Exits networks propose to evaluate the confidence of the model in the prediction of output at different stages within the network, with the option of exiting the computation path earlier, if the model is confident enough on the prediction. This has a twofold purpose: it allows to optimize the computational cost of the model tailored to the input difficulty, and moreover, it avoids the phenomenon of overthinking, a circumstance in which a correct prediction in the early stages of a network gets corrupted by noise in deeper layers (Kaya et al., 2019). The confidence evaluation of a model is carried out by attaching a number of auxiliary classifiers to the intermediate blocks of the network. Section 2.1.2 further illustrates different methods for the determination of the number and position of the exit points, and how to properly construct and evaluate the prediction confidence score of a network in the early stages. Before exploring the major contributions in the field, we present here a summary of the formulation of the Early Exits models presented in Scardapane et al. (2020b) and Scardapane et al. (2024), to which we refer for more details.
@@ A Neural Network can be seen as a composition of $b$ sequential blocks:
We proceed now to illustrate the major contributions in the field of Early Exits models. Section 2.1.1 presents the key architectures that have been introduced in the literature, Section 2.1.2 discusses papers that try to improve specific aspects of the learning process of these types of networks, while Section 2.1.3 focuses on applications to other tasks rather than classification. It should be noted that numerical results for the presented publications are not provided in this section. The domain of Early Exits lacks a standardized experimental protocol. We report on studies whose findings are based on a diverse array of datasets and evaluation architectures. A direct and fair quantitative analysis infeasible without extensive re-implementation. This heterogeneity highlights a critical gap in the current literature: the necessity for a structured benchmark. We argue that such a benchmark is essential to consistently compare different Dynamic Neural Network models, specifically accounting for inference speed and efficiency, the variability of computations in relation to input complexity, and performance quality in edge cases, such as small objects or targets located in the peripheral regions of the image. A further discussion on the topic is presented in the Discussion (Section 4).
@@ 2.1.1 Architectures
Following the principle that only a small fraction of inputs require the full computational effort of a network, while a large majority of the data samples can be classified correctly even in the very early stages of a network, Panda et al. (2016) is the first paper to inject Early Exits classifiers into a pretrained network. It takes a trained network and attaches linear classifiers at each layer, to establish if the layer is good enough to produce a correct classification output with confidence. The paper also includes a study on how to determine a good threshold to stop the computations.
BranchyNet (Teerapittayanon et al., 2016) improves on top of Panda et al. (2016) by presenting an end-to-end training procedure of the Neural Network: instead of having classifiers attached to the already trained network, this solution proposes to train the whole network end-to-end, with exit points composed of MLP branches for classification along the architecture. The network is trained by optimizing the sum of the loss functions associated with the classification at each exit point. The model uses an entropy score of a classifier’s output (see equation (4)) as a measurement of the model’s confidence in having a correct prediction. The contribution of each loss to the overall loss in (Teerapittayanon et al., 2016) is set manually, and the authors observe that weighting more early branches encourages a more discriminative feature learning in the early layers of the network, allowing more samples to exit with high confidence. A simplified schematic of the architecture can be seen in Figure 1(a). Other methods for computing the confidence score in different ways are discussed in more detail in Section 2.1.2. Kaya et al. (2019) proposes a slight variation with respect to the original BranchyNet paper, where the pre-trained weights are frozen and only the classifiers of the Early Exits are trained. Jie et al. (2021) adopts a CNN-based architecture that allows each sample to adaptively select its own optimal exit within a specific time budget. At each exit, there is a module trained through deep Reinforcement Learning which predicts the optimal exit for each sample.
@@
Tang et al. (2023) proposes an alternative method to perform Early Exits of tokens in the ViT for semantic segmentation. After each Transformer block in the encoder, an auxiliary head determines which of the tokens can be halted and passed through the decoder, and which need more computations. The paper further discusses the case where all the tokens of a given easy category get stopped in the early stages, losing the possibility to provide contextual information in the attention operation to the tokens of other classes. To avoid this, for each category, at least one token and at most top-$k$ of them are kept. This paper presents an approach that is in-between Token Skimming and Early Exits, which possible only due to the repetitive nature of the Tranformer architecture.
@@ 2.1.2 Methods
The exit policy is the logical component within an Early Exits network which is in charge of determining whether the model has processed the input enough to provide an intermediate output. Scardapane et al. (2020a) proposes a method to optimize the Early Exits decision jointly during training by adding a soft gate at each exit that decides a weighting strategy. The idea is to learn a weighting scheme to combine the output of the current prediction with the (supposedly improved) output of the next exit’s prediction. Essentially, the key idea is for the network to self-weight how much it should rely on the output of the current exit with respect to the output of the next Early Exit. This strategy allows training the model by exploiting a cross-entropy loss with respect to the final output solely; thus, it does not need to create a loss that is a weighted combination of all the losses computed on the intermediate exits. During inference, the added auxiliary gating value can be considered a standard binary classifier and can be thresholded to evaluate the probability of exit. Furthermore, the paper proposes a regularization term for the final loss that induces the training process to take into account also the average computational cost of the network. Similarly, in Chen et al. (2020) the same idea of learning the stopping policy for Early Exits together with the network parameters is considered. In this case, the solution views the stopping time as a latent variable conditioned on the input, interpreting the problem as a variational Bayes one. The training procedure to achieve this consists of decomposing the problem in two stages, necessary for learning both the predictive model and the stopping policy. In the first stage the model parameters are learned, assuming that the stop time always follows the best stopping distribution. In the second stage, the exit policy is learned by minimizing the reverse Kullback–Leibler divergence through a Reinforcement Learning algorithm.
@@
Pomponi et al. (2022) proposes a method to train jointly all the branches of a multi-exit model by weighting the predictions from each branch with a trained confidence score. Each confidence score is an approximation of the actual confidence score produced by the branch, which is calculated and regularized while training the rest of the model. The key idea is to allow the network to select the best Early Exits for a given input by associating to each auxiliary classifier an additional confidence score of exiting. The paper contribution consists in reinterpreting the output of the network probabilistically adding uncertainty quantification during training. Each confidence score is used as a parameter for a continuous relaxation of the Bernoulli distribution. Based on a similar idea, EERO (Valade et al., 2024) is a methodology that revisits the problem of Early Exiting using multiple classifiers with a reject option to better select the exiting head for each instance. The problem gets formalized as a classification with a reject option and establishes the optimal rule for a two-head scenario, deriving a data-driven procedure to define the rejection threshold. The probabilities of exiting at the different heads are calibrated using aggregation with exponential weights, with the aim of guaranteeing a fixed budget. Translated into the classification with reject option vocabulary, the question becomes whether the classifier should classify the instance or reject it.
Furthremore, Kim and Lee (2024) proposes a method based on Markov Decision Process (MDP) for the early-exit decision criterion. The key intuition is to use the sequentiality of the Early Exits decisions to model the exiting strategy as a Markov decision problem. A sequential inference process of each data sample can be represented as a single episode with a finite horizon. A Reinforcement Learning approach is then used to solve the problem by maximizing the accuracy as much as possible while minimizing computational cost. This solution allows the early-exit decisions to be dynamic instead of simply using a fixed early-exit criterion.
@@
Loss weighting methods try to determine the correct contribution of each exit output to the overall loss, with the aim of improving the learning of the network. It consists in determining the $\alpha_{i}$ values in equation (3). Wang et al. (2019a) proposes DynExit, a dynamic early-exit strategy for ResNet (He et al., 2016). This strategy introduces a dynamic loss-weight adjustment mechanism contributing to the weighting of the final loss as a weighted sum of all the losses, each computed at the respective exit. Instead of using manually tuned weighting values, in DynExit the weights become a trainable parameter with a dynamic range, which enables the loss-weight of different exit points. For this reason, the correct loss combination can be learned with all the other parameters, thus pondering the overall contribution of single exits to the learning process and increasing the overall accuracy. Furthermore, some hardware-aware optimizations of the implementation of the softmax and the cross entropy loss are presented, which are adapted to achieve high speed with low hardware complexity.
@@
In this paragraph, we cover those techniques that focus on the determination of the optimal amount and position of the exit points within an Early Exits network. The positioning of the exits is a straightforward problem where in most cases the default option is to have the exits equally distributed within the depth of the network. Nonetheless, some studies have researched the optimal position of these exit points. Laskaridis et al. (2020b) presents HAPI, a methodology for creating an optimal Early Exits network design by optimizing the placement of the intermediate exits as well as the number of exits within a network. The Early Exits variants of the network are represented in a graph-matrix form that allows the design to be solved as a mathematical optimization problem. The method changes the number and position of Early Exits along the network by adding and removing nodes in the network graph.
@@
Here we group all those methods that facilitate the learning of the objective of the Neural Network, taking also into account the Early Exits design of the architecture. Li et al. (2019) introduces two improvements to the training methods of Early Exits models. The first is an algorithm that tackles the problem of learning from different classifiers: the idea is to rescale the magnitude of gradients along the backpropagation path, allowing them to have a constant scale across the network for better stability. The second contribution is a knowledge transfer in combination with a self-distillation technique: a knowledge transfer path is applied among two adjacent classifiers, which can be considered as a residual connection of the classification. Concerning the self-distillation part, all the intermediate classifiers are also supervised with respect to the output of the last classifier. Other techniques of self-distillation will be discussed later in the eponymous paragraph.
@@
The train-test mismatch is a problem where all the prediction heads are optimized on all types of data in the training phase, while during inference the Early Exits of data will make the deeper layers of a network only see difficult inputs. The solution is to allow classifiers to focus more on hard samples to better learn how to specialize only to those cases. Han et al. (2022) proposes a method that allows to bridge this gap by imposing sample-wise weights on the loss of multiple exits. The key idea of the method is that the optimization objective should encourage each exit to emphasize different training samples by weighting their loss. To learn the appropriate weights, it leverages a weight prediction network (WPN). The WPN takes the training loss from all exits as input, producing the weights imposed on the samples at every exit. The backbone model and the WPN are optimized in a meta-learning manner. The meta objective for each classifier is then defined as the loss only on the samples that exit at a given classifier. Yu et al. (2022) proposes an alternative solution to the train-test mismatch problem. The solution starts from the assumption that a model is a weighted sum of weak prediction models and uses gradient boosting algorithms to minimize the loss function by iteratively training each weak model with gradient descent. In each training step, the model that points to the negative gradient of the loss function is selected as a new weak model, allowing each weak model to be considered supplementary to the ensemble of its previous weak models.
@@
Out-of-distribution (OOD) refers to a model’s ability to identify inputs that significantly differ from the distribution of the data the model was trained on. Xia and Bouganis (2023) explores tasks related to uncertainty estimation when it comes to Early Exits models. The proposed idea is to create an exit policy that instead of having a single threshold, uses a window-based approach. The intuition is that many downstream tasks for uncertainty estimation can be formulated as binary classification, thus only samples in a window near the binary decision boundary should be evaluated by later stages of a network. All the other samples outside of this window should be considered as OOD samples which can not be predicted and the network can exit earlier in order to avoid unnecessary computational cost. Also Meronen et al. (2024) proposes a Bayesian-based method to quantify and account for samples that are overconfidently wrongly estimated, trying to distinguish between samples where the model is uncertain and samples where the model is confidently wrong through probabilistic modeling. The motivation for improving uncertainty estimation is to allow intermediate classifiers to more accurately estimate the uncertainty in their prediction and thus better distinguish the cases where the overconfidence of the model results in an incorrect prediction in the early layers.
@@
Paper Main contribution Arch. Exit policy Application Section Wang and Li (2021) Dense knowledge distillation for each exit from all the later exits CNN Confidence Image classification 2.1.2 Wang et al. (2021b) Image elaboration at different scales with Early Exits Transf. Confidence Image classification 2.1.1 Wołczyk et al. (2021) Reuse the output of internal classifiers as inductive prediction CNN Confidence Image classification 2.1.2 Bakhtiarnia et al. (2022) Early exits audiovisual crowd counting Transf. - Image classification 2.1.3 Han et al. (2022) The loss is provided by a weighting network CNN Confidence Image classification 2.1.2 Kouris et al. (2022) Framework for Early Exit semantic segmentation CNN Confidence Image segmentation 2.1.3 Pomponi et al. (2022) Train by weighting the prediction with a trained confidence score CNN Confidence Image classification 2.1.2 Sabet et al. (2022) Temporal Early Exits for video object detection CNN Entropy Object detection 2.1.3 Sun et al. (2022) Weighted policy to alleviate gradient conflict problems CNN Confidence Image classification 2.1.2 Yu et al. (2022) Proposes a solution to the train-test mismatch problem CNN Confidence Image classification 2.1.2 Zhang et al. (2022) Experiments with various self-distillation techniques CNN Confidence Image classification 2.1.2 Han et al. (2023) Decouples the features extraction and the early classification CNN Confidence Image classification 2.1.1 Li et al. (2023a) learned component to decide where to place the classifiers CNN - Image classification 2.1.2 Li et al. (2023b) Early exit architecture applies to a Spiking Neural Network Spiking NN Confidence Image classification 2.1.3 Liu and Moon (2023) Balance the loss contribution by weight prediction CNN Confidence Image classification 2.1.2 Polina et al. (2023) Early Exits applied to Generative Adversarial Networks GAN quality threshold Image generation 2.1.3 Tang et al. (2023) Token Early Exit for semantic segmentation Transf. - Image segmentation 2.1.1 Xia and Bouganis (2023) Study of uncertainty estimation when it comes to Early Exit CNN window of confidence Image classification 2.1.2 Xu et al. (2023) Self-distillation to train Early Exit ViT models Transf. confidence threshold Image classification 2.1.2 Xue et al. (2023) ViT which has also the ability to read and store tokens Transf. - Image classification 2.1.1 Tang et al. (2023) Token Early Exit for semantic segmentation Transf. - Image segmentation 2.1.1 Bajpai and Hanawal (2024) ViT with Early Exits for image captioning Transf. Learned policy Image captioning 2.1.3

### v4

@@ Abstract
Model compression is essential in the deployment of large Computer Vision models on embedded devices. However, static optimization techniques (e.g. pruning, quantization, etc.) neglect the fact that different inputs have different complexities, thus requiring different amounts of computations. Dynamic Neural Networks allow conditioning the number of computations to the specific input. The current literature on the topic is very extensive and fragmented. We present a comprehensive survey that synthesizes and unifies existing Dynamic Neural Networks research in the context of Computer Vision. Additionally, we provide a logical taxonomy based on which component of the network is adaptive: the output, the computation graph or the input. Furthermore, we argue that Dynamic Neural Networks are particularly beneficial in the context of Sensor Fusion for better adaptivity, noise reduction and information prioritization. We present preliminary works in this direction. We complement this survey with a curated repository listing all the surveyed papers, each with a brief summary of the solution and the code base when available: https://github.com/DTU-PAS/awesome-dynn-for-cv.
@@ 1 Introduction
Furthermore, we argue that the adaptability of Dynamic Neural Networks is a particularly beneficial property in the context of Sensor Fusion—however, this is currently an underexplored research area. Studies suggest that, in the human brain, distinct parts are highly specialized in different tasks (Kandel et al., ), e.g. spatial information and object information are processed separately (Goodale and Milner, 1992). Furthermore, studies on different living organisms have shown how the loss of a sense gets compensated for by remaining sensory modalities (Rabinowitch et al., 2016). In Computer Vision, Multi-modal Sensor Fusion can be defined as the aptitude of a model to combine, interpret, and reason on information coming from different sensors and express various aspects of the same environment, aiming at a better perception of the surroundings in different contexts and conditions (Fung et al., 2017).
We observe that Sensor Fusion models can strongly benefit from the use of adaptive techniques for multiple reasons. Firstly, they allow sample-aware model efficiency. Furthermore, the learned adaptivity produces robustness against the noise generated by non-optimal environmental conditions in which the sensors are exposed. Finally, there is an intrinsic ability of these techniques to identify critical information by design. This allows organizing the processing according to input importance and computation budget constraints. As can be seen in Figure 2, Dynamic Neural Networks are gaining traction, with the number of papers steadily increasing year by year. While—as will be discussed later in this section—existing survey papers have covered similar directions, there is still a lack of a survey paper and a clear taxonomy that focuses on Computer Vision.
Figure 2: Overview of the publications considered in this survey, grouped by year and topic. In total, 163 publications have been reviewed: 65 for the Early Exits Section, 51 for the Dynamic Routing Section, 31 for the Token Skimming Section, and 16 in the Dynamic Sensor Fusion Section.
@@
    We propose a standardized evaluation template to address the critical lack of unified benchmarking protocols in the current literature, which aims at providing a fair and rigorous cross-methodology comparison methodology for future publications.
  • 5.
@@ 2.1 Early exits
The idea underlying of the development of Early Exits networks is rooted in the assumption that samples in a dataset can vary a lot in complexity. Following this principle, traditional Neural Networks tend to be overparametrized for most of the samples, which could potentially be processed with just a fraction of the parameters used. Furthermore, in the case of easy samples, Convolutional Neural Networks (CNNs) have been shown to be capable of extracting features in the early layers that are already sufficient for a confident classification (Panda et al., 2016). For these reasons, Early Exits networks propose to evaluate the confidence of the model in the prediction of output at different stages within the network, with the option of exiting the computation path earlier, if the model is confident enough in the prediction. This has a twofold purpose: it allows to optimize the computational cost of the model tailored to the input difficulty, and moreover, it avoids the phenomenon of overthinking, a circumstance in which a correct prediction in the early stages of a network gets corrupted by noise in deeper layers (Kaya et al., 2019). The confidence evaluation of a model is carried out by attaching a number of auxiliary classifiers to the intermediate blocks of the network. Section 2.1.2 further illustrates different methods for the determination of the number and position of the exit points, and how to properly construct and evaluate the prediction confidence score of a network in the early stages. Before exploring the major contributions in the field, we present here a summary of the formulation of the Early Exits models presented in Scardapane et al. (2020b) and Scardapane et al. (2024), to which we refer for more details.
@@ A Neural Network can be seen as a composition of $b$ sequential blocks:
We proceed now to illustrate the major contributions in the field of Early Exits models. Section 2.1.1 presents the key architectures that have been introduced in the literature, Section 2.1.2 discusses papers that try to improve specific aspects of the learning process of these types of networks, while Section 2.1.3 focuses on applications to other tasks rather than classification.
Although this survey considers Dynamic Neural Networks only in the context of Computer Vision, the theoretical formulation of Early Exits is not restricted to visual tasks. The idea at the base of these models—that different inputs have different complexities—applies equally to other fields, such as Natural Language Processing (NLP). Consequently, the methods proposed to compute the confidence score, like the entropy calculation of a classifier’s output, are independent of the data modality. Furthermore, solutions designed to optimize the exit policy or bridge the train-test mismatch gap present direct cross-domain applicability. For example, formulating the sequential early-exit decision criterion as a Markov Decision Process, or treating the stopping time as a latent variable within a variational Bayes problem, relies entirely on mathematical constructs independent of the input modality.
It should be noted that numerical results for the presented publications are not provided in this section. The domain of Early Exits lacks a standardized experimental protocol. We report on studies whose findings are based on a diverse array of datasets and evaluation architectures. A direct and fair quantitative analysis infeasible without extensive re-implementation. This heterogeneity highlights a critical gap in the current literature: the necessity for a structured benchmark. We argue that such a benchmark is essential to consistently compare different Dynamic Neural Network models, specifically accounting for inference speed and efficiency, the variability of computations in relation to input complexity, and performance quality in edge cases, such as small objects or targets located in the peripheral regions of the image. A further discussion on the topic is presented in the Discussion (Section 4).
@@ 2.1.1 Architectures
Following the principle that only a small fraction of inputs require the full computational effort of a network, while a large majority of the data samples can be classified correctly even in the very early stages of a network, Panda et al. (2016) is the first paper to inject Early Exits classifiers into a pretrained network. It takes a trained network and attaches linear classifiers at each layer, to establish whether the layer is good enough to produce a correct classification output with confidence. The paper also includes a study on how to determine a good threshold to stop the computations.
BranchyNet (Teerapittayanon et al., 2016) improves on top of Panda et al. (2016) by presenting an end-to-end training procedure for the Neural Network: instead of having classifiers attached to the already trained network, this solution proposes to train the whole network end-to-end, with exit points composed of MLP branches for classification along the architecture. The network is trained by optimizing the sum of the loss functions associated with the classification at each exit point. The model uses an entropy score of a classifier’s output (see equation (4)) as a measurement of the model’s confidence in having a correct prediction. The contribution of each loss to the overall loss in (Teerapittayanon et al., 2016) is set manually, and the authors observe that weighting more early branches encourages a more discriminative feature learning in the early layers of the network, allowing more samples to exit with high confidence. A simplified schematic of the architecture can be seen in Figure 1(a). Other methods for computing the confidence score in different ways are discussed in more detail in Section 2.1.2. Kaya et al. (2019) proposes a slight variation with respect to the original BranchyNet paper, where the pre-trained weights are frozen and only the Early Exits classifiers are trained. Jie et al. (2021) adopts a CNN-based architecture that allows each sample to adaptively select its own optimal exit within a specific time budget. At each exit, there is a module trained through deep Reinforcement Learning which predicts the optimal exit for each sample.
@@
Tang et al. (2023) proposes an alternative method to perform Early Exits of tokens in the ViT for semantic segmentation. After each Transformer block in the encoder, an auxiliary head determines which of the tokens can be halted and passed through the decoder, and which need more computations. The paper further discusses the case where all the tokens of a given easy category get stopped in the early stages, losing the possibility to provide contextual information in the attention operation to the tokens of other classes. To avoid this, for each category, at least one token and at most top-$k$ of them are kept. This paper presents an approach that is in-between Token Skimming and Early Exits, which is possible only due to the repetitive nature of the Tranformer architecture.
@@ 2.1.2 Methods
The exit policy is the logical component within an Early Exits network which is in charge of determining whether the model has processed the input enough to provide an intermediate output. Scardapane et al. (2020a) proposes a method to optimize the Early Exits decision jointly during training by adding a soft gate at each exit that decides a weighting strategy. The idea is to learn a weighting scheme to combine the output of the current prediction with the (supposedly improved) output of the next exit’s prediction. Essentially, the key idea is for the network to self-weight how much it should rely on the output of the current exit with respect to the output of the next Early Exit. This strategy allows training the model by exploiting a cross-entropy loss with respect to the final output solely; thus, it does not need to create a loss that is a weighted combination of all the losses computed on the intermediate exits. During inference, the added auxiliary gating module can be considered a standard binary classifier and can be thresholded to evaluate the probability of exit. Furthermore, the paper proposes a regularization term for the final loss that induces the training process to take into account also the average computational cost of the network. Similarly, in Chen et al. (2020) the same idea of learning the stopping policy for Early Exits together with the network parameters is considered. In this case, the solution views the stopping time as a latent variable conditioned on the input, interpreting the problem as a variational Bayes one. The training procedure to achieve this consists of decomposing the problem in two stages, necessary for learning both the predictive model and the stopping policy. In the first stage the model parameters are learned, assuming that the stop time always follows the best stopping distribution. In the second stage, the exit policy is learned by minimizing the reverse Kullback–Leibler divergence through a Reinforcement Learning algorithm.
@@
Pomponi et al. (2022) proposes a method to train jointly all the branches of a multi-exit model by weighting the predictions from each branch with a trained confidence score. Each confidence score is an approximation of the actual confidence score produced by the branch, which is calculated and regularized while training the rest of the model. The key idea is to allow the network to select the best Early Exits for a given input by associating to each auxiliary classifier an additional confidence score of exiting. The paper contribution consists in reinterpreting the output of the network probabilistically adding uncertainty quantification during training. Each confidence score is used as a parameter for a continuous relaxation of the Bernoulli distribution. Based on a similar idea, EERO (Valade et al., 2024) is a methodology that revisits the problem of Early Exiting using multiple classifiers with a reject option to better select the exiting head for each instance. The problem is formalized as a classification with a reject option and establishes the optimal rule for a two-head scenario, deriving a data-driven procedure to define the rejection threshold. The probabilities of exiting at the different heads are calibrated using aggregation with exponential weights, with the aim of guaranteeing a fixed budget. Translated into the classification with reject option vocabulary, the question becomes whether the classifier should classify the instance or reject it.
Furthermore, Kim and Lee (2024) proposes a method based on Markov Decision Process (MDP) for the early-exit decision criterion. The key intuition is to use the sequentiality of the Early Exits decisions to model the exiting strategy as a Markov decision problem. A sequential inference process of each data sample can be represented as a single episode with a finite horizon. A Reinforcement Learning approach is then used to solve the problem by maximizing the accuracy as much as possible while minimizing computational cost. This solution allows the early-exit decisions to be dynamic instead of simply using a fixed early-exit criterion.
@@
Loss weighting methods try to determine the correct contribution of each exit output to the overall loss, with the aim of improving the learning of the network. It consists in determining the $\alpha_{i}$ values in equation (3). Wang et al. (2019a) proposes DynExit, a dynamic early-exit strategy for ResNet (He et al., 2016). This strategy introduces a dynamic loss-weight adjustment mechanism contributing to the weighting of the final loss as a weighted sum of all the losses, each computed at the respective exit. Instead of using manually tuned weighting values, in DynExit the weights become a trainable parameter with a dynamic range, which enables the loss-weight of different exit points. For this reason, the correct loss combination can be learned with all the other parameters, thus weighting the overall contribution of single exits to the learning process and increasing the overall accuracy. Furthermore, some hardware-aware optimizations of the implementation of the softmax and the cross entropy loss are presented, which are adapted to achieve high speed with low hardware complexity.
@@
In this paragraph, we cover those techniques that focus on the determination of the optimal number and position of the exit points within an Early Exits network. The positioning of the exits is a straightforward problem where in most cases the default option is to have the exits equally distributed within the depth of the network. Nonetheless, some studies have researched the optimal position of these exit points. Laskaridis et al. (2020b) presents HAPI, a methodology for creating an optimal Early Exits network design by optimizing the placement of the intermediate exits as well as the number of exits within a network. The Early Exits variants of the network are represented in a graph-matrix form that allows the design to be solved as a mathematical optimization problem. The method changes the number and position of Early Exits along the network by adding and removing nodes in the network graph.
@@
Here we group all those methods that facilitate the learning of the objective of the Neural Network, taking also into account the Early Exits design of the architecture. Li et al. (2019) introduces two improvements to the training methods of Early Exits models. The first is an algorithm that tackles the problem of learning from different classifiers: the idea is to rescale the magnitude of gradients along the backpropagation path, allowing them to have a constant scale across the network for better stability. The second contribution is a knowledge transfer in combination with a self-distillation technique: a knowledge transfer path is applied among two adjacent classifiers, which can be considered as a residual connection of the classification. Concerning the self-distillation part, all the intermediate classifiers are also supervised using the output of the last classifier. Other techniques of self-distillation will be discussed later in the eponymous paragraph.
@@
The train-test mismatch is a problem where all the prediction heads are optimized on all types of data in the training phase, while during inference the early exiting of data will make the deeper layers of a network only see difficult inputs. The solution is to allow classifiers to focus more on hard samples to better learn how to specialize only to those cases. Han et al. (2022) proposes a method that allows to bridge this gap by imposing sample-wise weights on the loss of multiple exits. The key idea of the method is that the optimization objective should encourage each exit to emphasize different training samples by weighting their loss. To learn the appropriate weights, it leverages a weight prediction network (WPN). The WPN takes the training loss from all exits as input, producing the weights imposed on the samples at every exit. The backbone model and the WPN are optimized in a meta-learning manner. The meta objective for each classifier is then defined as the loss only on the samples that exit at a given classifier. Yu et al. (2022) proposes an alternative solution to the train-test mismatch problem. The solution starts from the assumption that a model is a weighted sum of weak prediction models and uses gradient boosting algorithms to minimize the loss function by iteratively training each weak model with gradient descent. In each training step, the model that points to the negative gradient of the loss function is selected as a new weak model, allowing each weak model to be considered supplementary to the ensemble of its previous weak models.
@@
Out-of-distribution (OOD) detection refers to a model’s ability to identify inputs that significantly differ from the distribution of the data the model was trained on. Xia and Bouganis (2023) explores tasks related to uncertainty estimation when it comes to Early Exits models. The proposed idea is to create an exit policy that instead of having a single threshold, uses a window-based approach. The intuition is that many downstream tasks for uncertainty estimation can be formulated as binary classification, thus only samples in a window near the binary decision boundary should be evaluated by later stages of a network. All the other samples outside of this window should be considered as OOD samples which can not be predicted and the network can exit earlier in order to avoid unnecessary computational cost. Also Meronen et al. (2024) proposes a Bayesian-based method to quantify and account for samples that are overconfidently wrongly estimated, trying to distinguish between samples where the model is uncertain and samples where the model is confidently wrong through probabilistic modeling. The motivation for improving uncertainty estimation is to allow intermediate classifiers to more accurately estimate the uncertainty in their prediction and thus better distinguish the cases where the overconfidence of the model results in an incorrect prediction in the early layers.
@@
Paper Main contribution Arch. Exit policy Application Section Wang and Li (2021) Dense knowledge distillation for each exit from all the later exits CNN Confidence Image classification 2.1.2 Wang et al. (2021b) Image elaboration at different scales with Early Exits Transf. Confidence Image classification 2.1.1 Wołczyk et al. (2021) Reuse the output of internal classifiers as inductive prediction CNN Confidence Image classification 2.1.2 Bakhtiarnia et al. (2022) Early exits audiovisual crowd counting Transf. - Image classification 2.1.3 Han et al. (2022) The loss is provided by a weighting network CNN Confidence Image classification 2.1.2 Kouris et al. (2022) Framework for Early Exit semantic segmentation CNN Confidence Image segmentation 2.1.3 Pomponi et al. (2022) Train by weighting the prediction with a trained confidence score CNN Confidence Image classification 2.1.2 Sabet et al. (2022) Temporal Early Exits for video object detection CNN Entropy Object detection 2.1.3 Sun et al. (2022) Weighted policy to alleviate gradient conflict problems CNN Confidence Image classification 2.1.2 Yu et al. (2022) Proposes a solution to the train-test mismatch problem CNN Confidence Image classification 2.1.2 Zhang et al. (2022) Experiments with various self-distillation techniques CNN Confidence Image classification 2.1.2 Han et al. (2023) Decouples the features extraction and the early classification CNN Confidence Image classification 2.1.1 Li et al. (2023a) learned component to decide where to place the classifiers CNN - Image classification 2.1.2 Li et al. (2023b) Early exit architecture applies to a Spiking Neural Network Spiking NN Confidence Image classification 2.1.3 Liu and Moon (2023) Balance the loss contribution by weight prediction CNN Confidence Image classification 2.1.2 Polina et al. (2023) Early Exits applied to Generative Adversarial Networks GAN quality threshold Image generation 2.1.3 Tang et al. (2023) Token Early Exit for semantic segmentation Transf. - Image segmentation 2.1.1 Xia and Bouganis (2023) Study of uncertainty estimation when it comes to Early Exit CNN window of confidence Image classification 2.1.2 Xu et al. (2023) Self-distillation to train Early Exit ViT models Transf. confidence threshold Image classification 2.1.2 Xue et al. (2023) ViT which has also the ability to read and store tokens Transf. - Image classification 2.1.1 Bajpai and Hanawal (2024) ViT with Early Exits for image captioning Transf. Learned policy Image captioning 2.1.3